# BadSFL: Backdoor Attack against Scaffold Federated Learning

### v2

@@ Abstract
Federated learning (FL) enables the training of deep learning models on distributed clients to preserve data privacy. However, this learning paradigm is vulnerable to backdoor attacks, where malicious clients can upload poisoned local models to embed backdoors into the global model, leading to attacker-desired predictions. Existing backdoor attacks mainly focus on FL with independently and identically distributed (IID) scenarios, while real-world FL training data are typically non-IID. Current strategies for non-IID backdoor attacks suffer from limitations in maintaining effectiveness and durability. To address these challenges, we propose a novel backdoor attack method, BadSFL, specifically designed for the FL framework using the scaffold aggregation algorithm in non-IID settings. BadSFL leverages a Generative Adversarial Network (GAN) based on the global model to complement the training set, achieving high accuracy on both backdoor and benign samples. It utilizes a specific feature as the backdoor trigger to ensure stealthiness, and exploits the Scaffold’s control variate to predict the global model’s convergence direction, ensuring the backdoor’s persistence. Extensive experiments on three benchmark datasets demonstrate the high effectiveness, stealthiness, and durability of BadSFL. Notably, our attack remains effective over 60 rounds in the global model and up to 3 times longer than existing baseline attacks after stopping the injection of malicious updates.
Introduction
@@
The distribution of training data across clients plays a crucial role in the performance of Federated Learning (FL) systems. Two scenarios typically arise: 1) the training data is IID among clients; 2) A more realistic scenario where the data is non-IID, meaning the data characteristics vary significantly across clients. For IID scenarios, FedAvg (McMahan et al. 2017) stands out as the leading FL method, setting the standard for server-side model updates by aggregating model parameters from clients effectively. However, its performance deteriorates in non-IID scenarios, where data heterogeneity causes update drifts from individual clients, ultimately degrading convergence. (Li et al. 2019). To address this challenge, Scaffold (Karimireddy et al. 2020) has been proposed as a promising solution, using a control variate to correct the drift induced by the data heterogeneity.
However, while FL’s distributed training paradigm allows clients to have full control over their data for privacy protection, it also introduces a non-negligible security threat: backdoor attacks, where a malicious client can manipulate the local training process to embed targeted misbehaviors into the global model through poisoned updates (Chen et al. 2017).
Existing federated learning backdoor attacks (Bagdasaryan et al. 2020) (Xie et al. 2019) (Sun et al. 2019) (Wang et al. 2020a) (Dai and Li 2023) mainly focus on IID scenarios, where attackers compromise clients with full knowledge of the dataset distribution to craft highly effective attacks. Unfortunately, these methods have not demonstrated their feasibility and effectiveness in non-IID scenarios. Specifically, attackers lacking full knowledge of the dataset distribution often result in training backdoor models diverging significantly from the global model, severely compromising overall model performance. Consequently, directly applying attack methods from IID scenarios to non-IID scenarios is ineffective. Some efforts have performed backdoor attacks on non-IID scenarios (Alam, Sarkar, and Maniatakos 2023) (Jeong, Hwang, and Chung 2021) (Naseri, Han, and De Cristofaro 2023) (Ye et al. 2024). However, their effectiveness remains limited against advanced defenses like Scaffold, since they do not take the control variate into consideration.
@@
In this paper, we propose BadSFL, a novel backdoor attack specifically targeting Scaffold Federated Learning (SFL), to successfully plant a backdoor function into a global model without catastrophically corrupt model performance on benign sample inference. Firstly, as the attacker only has partial knowledge of the dataset distribution in the FL system, we leverage a GAN to generate fake samples that belong to other clients to supplement the dataset, simulating a full knowledge of the dataset distribution. With a supplemented dataset for backdoor training, we get a backdoor model achieving high accuracy in both backdoor tasks and benign tasks. Secondly, we use a distinctive feature of a category as the backdoor trigger to maintain the stealthiness of BadSFL. Thirdly, we exploit the global control variate, a specific feature of SFL, as a reference to predict the global model’s convergence direction. This optimization enhances the durability of the embedded backdoor function within the global model. We evaluate the effectiveness of BadSFL on the MNIST, CIFAR-10, and CIFAR-100, demonstrating high accuracy on both backdoor and primary tasks. Moreover, the embedded backdoor function persists in the global model for over 60 rounds and lasts for 3 times longer than baseline attacks after the attacker stops injecting malicious updates. Finally, we show that BadSFL remains highly effective when using three defense methods simultaneously.
@@
  • We propose BadSFL, the first backdoor attack against scaffold federated learning on non-IID scenarios.
@@
  • We enhance the durability of the backdoor, ensuring it persists for over 60 rounds and lasts 3 times longer than baseline attacks.
@@
  • We conduct extensive experiments on three benchmark datasets. The results demonstrate that our method achieves high effectiveness.
Background and Related Work
Non-IID Scenarios in FL
@@
In FL, non-IID refers to significant differences in data distributions among clients (Kairouz et al. 2021) (Li et al. 2021b) (Mendieta et al. 2022). This discrepancy between local data distributions in NON-IID scenarios can lead to inconsistencies between the local optima and the global optima. This inconsistency results in a drift in local model updates, where local models move towards their own local optima which can be far from the global optima (Karimireddy et al. 2020). Consequently, even when averaging these local models may yield a global model far from the true global optima (Li et al. 2020) (Wang et al. 2020b) (Wang et al. 2020c) (Karimireddy et al. 2020) (Lee et al. 2022), especially with numerous local epochs. Figure 1 illustrates this issue: while global optima align with local optima in IID scenarios. non-IID can cause the global optima to be distant from individual local optima, which is known as the client-drift phenomenon, leading to slow and unstable convergence in the FL training process.
Figure 1: Model Avg under IID and non-IID scenarios.
@@
SCAFFOLD. Several FL algorithms have been proposed to address the above challenges, with SCAFFOLD (Karimireddy et al. 2020) being the most practical solution. It tackles the client-drift problem through the control variates (variance reduction techniques) for both the server and the clients. These control variates estimate the update direction of the global model and local client models and serve to correct local updates based on the drift, thereby mitigating the divergence between local and global optima. Therefore, we mainly focus on designing backdoor attacks targeting scaffold federated learning (SFL) in this paper.
Backdoor Attacks in FL
@@
Backdoor attacks pose a significant threat to the security of deep learning models, where malicious clients embed hidden triggers within models that cause misclassification during inference while maintaining normal performance on clean data (Li et al. 2022) (Gu et al. 2019). In FL, adversaries can deploy backdoor attacks by exploiting compromised clients to manipulate local updates, thereby generating poisoned models that corrupt the global model upon aggregation (Bagdasaryan et al. 2020)(Sun et al. 2019)(Wang et al. 2020a)(Xie et al. 2019)(Zhang et al. 2022) (Li, He, and Song 2021). Backdoor attacks typically involve a combination of model replacement and data poisoning. In model replacement, the adversary substitutes legitimate models with manipulated ones (Bagdasaryan et al. 2020), while data poisoning involves injecting poisoned data, containing the backdoor trigger, into the training datasets of compromised clients. To inject these triggers, the adversary can manipulate the dataset set by flipping data labels or adding a unique pixel pattern to images (Bagdasaryan et al. 2020)(Wang et al. 2020a)(Alam, Lamri, and Maniatakos 2023). Afterward, they strategically adjust training parameters and scale updates to optimize the impact of the attack while evading detection by the anomaly detector deployed at the central aggregation server (Blanchard et al. 2017)(Nguyen et al. 2022). In this paper, we mainly focus on data poisoning attacks.
Threat Model
@@
Attack scenarios. We consider an attacker who aims to inject backdoors into SFL to make the victim clients predict the desired wrong output over a triggered input. The attacker has partial knowledge of the full dataset in the training stage. Specifically, the attacker trains a local model with a backdoor trigger function and submits poisoned local updates along with the control variate to the server for Scaffold aggregation. During inference, the attack manipulates predictions to produce the attacker’s desired outputs when inputs from benign clients meet the trigger conditions.
@@
    Stealthiness: the backdoor should be robust against potential defenses.
@@
Attack capability. We assume the attacker can compromise at least one client, acting as a benign client during the training process, thereby allowing him to operate covertly within the system. Additionally, participation in FL provides the attacker with full knowledge of the model structure, facilitated by the consensus among all clients on a common learning objective. This enables the insertion of backdoor triggers, modification of sample labels, and manipulation of local training updates. It is essential to note that the attacker cannot control the server or directly manipulate the aggregation procedure or the global model. They also lack access to data and models from non-compromised clients.
Challenges with Backdoor Attacks in SFL
Performing a backdoor attack in SFL presents the following challenges.
Limited knowledge: In non-IID scenarios, a primary challenge arises from the attacker’s lack of knowledge of the dataset distribution across clients. Unlike IID scenarios where a centralized understanding of the dataset facilitated manipulation, non-IID scenarios involve decentralized and diverse data distributions. This results in three issues: (1) Direct backdoor strategies can cause significant performance degradation on benign samples, leading to the rejection of the global model. (2) The variability in data distributions increases the difference between local and global models, making malicious models more detectable. (3) Averaging poisoned models with the global model degrades its performance on the primary task, as shown in Figure 2, where accuracy drops significantly when a poisoned model is aggregated.
@@
Control variate: In SFL, control variate (denoted as $c_{i}$) is used to correct the client drift and align local models with the global model. If attackers strictly follow protocols and use the $c_{i}$) to correct their malicious models during the triggering planting process, the effectiveness of the attack can be reduced. Conversely, if an attacker chooses to manipulate the $c_{i}$ inappropriately, introducing a malicious $c$ to the server, it could lead to a potential corruption of the global model.
Backdoor catastrophic forgetting: Catastrophic forgetting (Kirkpatrick et al. 2017) occurs when neural networks forget previously learned tasks upon learning new ones. This can cause backdoors to lose effectiveness over time, as demonstrated by Wang et al. (Wang et al. 2020a). If attackers stop uploading malicious updates, the backdoor function may eventually be erased by benign updates. As shown in Figure 2, the accuracy of backdoor tasks declines sharply over time, with the backdoor function vanishing around round 65. Although various methods (Zhang et al. 2022)(Alam, Sarkar, and Maniatakos 2023)(Dai and Li 2023)(Wen et al. 2022) have been proposed to address this issue, none have proven effective in SFL.
Algorithm 1 BadSFL attack Algorithm
@@
Overview. To overcome the challenges, we propose BadSFL, which is detailed in Algorithm 1. BadSFL mainly consists of 4 steps: Step 1: Initialization. The attacker initiates the attack by downloading the global model $w_{g}$ and controlling the variate $c$ from the server. Subsequently, the attacker updates the local model $w_{p}$ and discriminator $D$ using the downloaded global model $w_{g}$. Step 2: GAN-based Training for Data Supplementation. The attacker performs GAN training on generator $G$ and discriminator $D$. The training terminates upon the convergence of the generator, signifying its ability to generate realistic fake samples in class $C$ that do not belong to $D^{i}$ but rather originate from other clients’ datasets. Then the generator $G$ is utilized to generate a number of samples in class $c$ forming in dataset $D_{f}$. This dataset $D_{f}$ is then merged with the attacker’s original dataset $D^{i}$ to create a new dataset $D_{c}$. Step 3: Backdoor Sample Selection and Trigger Injection. With dataset $D_{c}$, the attacker selects specific samples with a characteristic feature to serve as backdoor samples. These samples are then relabeled to a target class $x$ as the backdoor target class that is different from their original labels. The attacker organizes these manipulated samples into a separate dataset $D_{b}$ and merges it with $D_{c}$ to finalize the dataset $D_{p}$ for backdoor training. Step 4: Backdoor Model Training and Optimization. The attacker proceed to trains the local model $w_{p}$ based on the dataset $D_{p}$. During the training process, the attacker follows the equation 3 to optimize the backdoor objective. Upon convergence, the backdoor model update $\Delta w_{p}$ and the corresponding control variate $\Delta c_{p}$ are obtained and can be uploaded to the server.

### v3

# Mind the Cost of Scaffold! Benign Clients May Even Become Accomplices of Backdoor Attack

@@ Abstract
By using a control variate to calibrate the local gradient of each client, Scaffold has been widely known as a powerful solution to mitigate the impact of data heterogeneity in Federated Learning. Although Scaffold achieves significant performance improvements, we show that this superiority is at the cost of increased security vulnerabilities. Specifically, this paper presents BadSFL, the first backdoor attack targeting Scaffold, which turns benign clients into accomplices to amplify the attack effect. The core idea of BadSFL is to uniquely tamper with the control variate to subtly steer benign clients’ local gradient updates towards the attacker’s poisoned direction, effectively turning them into unwitting accomplices, significantly enhancing the backdoor persistence. Additionally, BadSFL leverages a GAN-enhanced poisoning strategy to enrich the attacker’s dataset, maintaining high accuracy on both benign and backdoored samples while remaining stealthy. Extensive experiments demonstrate that BadSFL achieves superior attack durability, maintaining effectiveness for over 60 global rounds—lasting up to three times longer than existing baselines even after ceasing malicious model injections.
1 Introduction
@@
Federated Learning (FL) enables distributed model training while preserving client data privacy. However, the effectiveness of FL models heavily depends on the distribution of training data across clients. Two scenarios typically arise: 1) IID data, where training data is uniformly distributed across clients, and 2) non-IID data, a more realistic setting where data characteristics vary significantly across clients. For IID scenarios, FedAvg [27] stands out as the leading FL method, setting the standard for server-side model updates by aggregating model parameters from clients. However, its performance deteriorates in non-IID scenarios, where data heterogeneity causes update drifts from individual clients, ultimately degrading convergence [22].
To address this challenge, Scaffold [12] was introduced as a robust FL method designed to mitigate client update drift through a correction mechanism based on control variates, thereby enhancing model convergence in non-IID settings. The control variate is essentially an estimate of the difference between a client’s local gradient and the global gradient, which helps align the local updates with the global objective. Scaffold reduces variance in the updates caused by data heterogeneity, making it particularly effective for scenarios where clients have diverse data distributions.
However, Scaffold Federated Learning (SFL) not only changes the way FL models converge but also affects their robustness against adversarial manipulations. In particular, malicious clients in FL can exploit model update mechanisms to introduce backdoor behaviors, embedding hidden misbehavior into the global model [6]. While backdoor attacks have been extensively studied in FL [3, 40, 34, 36, 7], most existing works focus on IID scenarios where attackers have full knowledge of the dataset distribution and can easily craft poisoned updates. In contrast, non-IID data distributions introduce additional constraints, making it harder for attackers to align poisoned models with the global model without significantly degrading overall performance. Although recent studies have explored backdoor attacks in non-IID FL [2, 10, 29, 42], they have largely overlooked the unique security implications introduced by SFL. The question this paper aims to address is: if the new mechanisms of SFL (i.e., control variate for update drift correction) can bring new security threats, and unintentionally facilitate backdoor attacks in non-IID settings?
Our answer to the above question is affirmative. Our new insight is that Scaffold’s reliance on control variates introduces a novel attack surface: its correction mechanism, designed to stabilize training by aligning local updates with the global objective, can inadvertently amplify the impact of malicious updates. More critically, this mechanism allows an attacker to influence the control variate itself, effectively co-opting benign clients to “assist in the mischief”. Since all clients use the control variate to adjust their local gradients during updates, a tampered variate can subtly steer these honest clients’ gradients toward the attacker’s poisoned direction. This amplifies the backdoor’s reach, making Scaffold more susceptible to sophisticated attacks than standard FL methods like FedAvg, which lack such a correction mechanism.
Figure 1: Model averaging under IID and non-IID scenarios.
@@
To exploit the above insight, we propose BadSFL, a novel backdoor attack specifically targeting Scaffold Federated Learning, to successfully implant a backdoor function into a global model without catastrophically corrupting model performance on benign sample inference. Unlike prior attacks, BadSFL leverages Scaffold’s correction dynamics to enhance both the stealth and durability of the backdoor, revealing a critical vulnerability in SFL methods. BadSFL operates as follows: Firstly, as the attacker only has partial knowledge of the dataset distribution in the FL system, he leverages a GAN to generate fake samples that belong to other clients to supplement the dataset, simulating a full knowledge of the dataset distribution. With the supplemented dataset for backdoor training, he gets a backdoor model achieving high accuracy in both backdoor tasks and benign tasks. Secondly, the attacker uses a distinctive feature of a category as the backdoor trigger to maintain the attack stealthiness. Thirdly, the attacker exploits the global control variate, as a reference to predict the global model’s convergence direction. This optimization significantly enhances the durability of the embedded backdoor function within the global model.
We evaluate BadSFL on the MNIST, CIFAR-10, and CIFAR-100 datasets, demonstrating its high accuracy on both backdoor and primary tasks. Moreover, the embedded backdoor function persists in the global model for over 60 rounds and lasts 3 times longer than baseline attacks after the attacker stops injecting malicious updates. Finally, we show that BadSFL remains highly effective when using four defense methods simultaneously.
@@
  • We propose BadSFL, the first backdoor attack against SFL on non-IID scenarios.
@@
  • We enhance the backdoor durability, ensuring it persists for over 60 rounds and lasts 3×3$\times$3 × longer than baselines.
@@
  • We conduct extensive experiments on three benchmark datasets, demonstrating high effectiveness of our attack.
2 Background and Related Work
2.1 Non-IID Scenarios in FL
@@
In FL, non-IID refers to significant differences in data distributions among clients [11, 20, 28]. This discrepancy between local data distributions in non-IID scenarios can lead to inconsistencies between the local optima and the global optima. This inconsistency results in a drift in local model updates, where local models move towards their own local optima which can be far from the global optima [12]. Consequently, averaging these local models may yield a global model far from the true global optima [21, 37, 38, 12, 16], especially with numerous local epochs. As shown in Figure 1, while the global optima aligns with the local optima in IID scenarios, non-IID can cause the global optima to be distant from individual local optima, which is known as the client-drift phenomenon, leading to slow and unstable convergence in the FL training process.
Algorithm 1 Scaffold Algorithm in Federated Learning
Sever Input: local datasets $D^{i}$, number of client $K$, number of communication rounds $R$, number of local epochs $E$
Client Input: local control variates $c_{i}$, local step-size $\eta_{l}$
Server Updates:
   $c^{t}\leftarrow 0$;
   for each round r = 1, …, $R$ do:
       randomly selected clients $S^{t}\subseteq\{1,\ldots,K\}$
       for $i\in S^{t}$ in parallel do:
          send $w^{t},c^{t}\rightarrow i$
          $\Delta w^{t}_{i},\Delta c^{t}_{i}\leftarrow$ Local Update($i,w^{t},c^{t}$)
       $w^{t}+1\leftarrow w^{t}-\eta\sum_{i\in S^{t}}\Delta w^{t}_{i}$
       $c^{t}+1\leftarrow c^{t}+\frac{1}{K}\Delta c$
   end for
Local Updates:
   Local client $i$ get $w^{t},c^{t}$ from server
   training model with $D_{i}$ get gradient $g_{i}(w_{i})$
   update local model:
       $w_{i}\leftarrow w_{i}-\eta_{l}*(g_{i}(w_{i})-c_{i}+c)$
   update local control variates:
       $c^{t+1}_{i}\leftarrow(i)\hskip 10.00002ptg_{i}(w_{i}),or\hskip 10.00002pt(ii)\hskip 10.00002ptc_{i}-c+\frac{1}{K*\eta_{l}}*(w^{t}-w_{i})$
   $\Delta w^{t}_{i}\leftarrow w_{i}-w^{t}$
   $\Delta c^{t}_{i}\leftarrow c^{t+1}_{i}-c_{i}$
   return ($\Delta w^{t}_{i},\Delta c^{t}_{i}$)
@@
Scaffold. Several FL algorithms have been proposed to address the above challenges, with Scaffold [12] being the most practical solution. It tackles the client-drift problem through the control variates (variance reduction techniques) for both the server and the clients. These control variates estimate the update direction of the global model and local client models and serve to correct local updates based on the drift, thereby mitigating the divergence between local and global optima (Alg. 1). In this paper, we mainly focus on designing backdoor attacks targeting SFL.
2.2 Backdoor Attacks against FL
@@
Backdoor attacks pose a significant threat to deep learning models, where malicious clients embed hidden triggers within models that cause misclassification during inference while maintaining normal performance on clean data [23]. In FL, adversaries can deploy backdoor attacks by exploiting compromised clients to manipulate local updates, thereby generating poisoned models that corrupt the global model upon aggregation [3, 34, 36, 40, 45, 19].
Backdoor attacks typically involve a combination of model replacement and data poisoning. In model replacement, the adversary substitutes legitimate models with manipulated ones [3], while data poisoning involves injecting poisoned data containing the backdoor trigger into the training datasets of compromised clients. To inject these triggers, the adversary can manipulate the dataset by flipping data labels or adding a unique pixel pattern to the training samples [3, 36, 1]. Afterward, he strategically adjusts training parameters and scales updates to optimize the impact of the attack while evading detection by the anomaly detector deployed at the central aggregation server [5, 30].
In this paper, we focus on data poisoning-based backdoor attacks. While most existing backdoor attacks target FL under IID scenarios [43, 9, 31, 40, 25], real-world FL deployments often involve non-IID distributions, posing additional challenges for effective backdoor injection. To the best of our knowledge, no prior research has specifically explored backdoor attacks against FL with the Scaffold aggregation algorithm. We bridge this gap by investigating a novel backdoor attack targeting SFL, leveraging its unique control variate mechanism to enhance the effectiveness, stealthiness, and persistence of the attack.
2.3 Threat Model
@@
Attack scenarios. We consider an attacker who aims to inject backdoors into SFL, make the final model predict the desired wrong output over a triggered input. The attacker has partial knowledge of the full dataset in the training stage. Specifically, the attacker trains a local model with a backdoor trigger function and submits poisoned local updates along with the control variate to the server for Scaffold aggregation. During inference, the attacker manipulates predictions to produce the attacker’s desired outputs when inputs meet the trigger conditions.
@@
    Robustness: the backdoor should be robust against potential defenses.
@@
Attacker’s capability. We assume the attacker can compromise at least one client during the training process, thereby allowing him to operate covertly within the system. Additionally, participation in FL provides the attacker with full knowledge of the model structure, facilitated by the consensus among all clients on a common learning objective. This enables the insertion of backdoor triggers, the modification of sample labels, and the manipulation of local training updates. It is essential to note that the attacker cannot control the server or directly manipulate the aggregation procedure or the global model. They also lack access to data and models from non-compromised clients.
@@
3 Challenges with Backdoor Attacks in SFL
Performing a backdoor attack in SFL presents the following challenges.  Limited knowledge. In non-IID scenarios, a primary challenge arises from the attacker’s lack of knowledge of the dataset distribution across clients. Unlike IID scenarios, where a centralized understanding of the dataset facilitates manipulation, non-IID scenarios involve decentralized and diverse data distributions. This results in three issues: (1) Direct backdoor strategies can cause significant performance degradation on benign samples, leading to the rejection of the global model; (2) The variability in data distributions increases the difference between local and global models, making malicious models more detectable; (3) Averaging poisoned models with the global model degrades its performance on the primary task, as shown in Figure 2, where accuracy drops significantly when a poisoned model is aggregated.  Control variate. In SFL, control variate (denoted as $c_{i}$) is used to correct the client drift and align local models with the global model. If attackers strictly follow protocols and use the $c_{i}$) to correct their malicious models during the triggering planting process, the effectiveness of the attack can be reduced. Conversely, if an attacker chooses to manipulate the $c_{i}$ inappropriately, introducing a malicious $c$ to the server, it could lead to a potential corruption of the global model.  Backdoor catastrophic forgetting. Catastrophic forgetting [13] occurs when neural networks forget previously learned tasks upon learning new ones. This can cause backdoors to lose effectiveness over time [36]. If attackers stop uploading malicious updates, the backdoor function may eventually be erased by benign updates. As shown in Figure 2, the accuracy of backdoor tasks declines sharply over time, with the backdoor function vanishing around round 65. Although various methods [45, 2, 7, 39] have been proposed to address this issue, none have proven effective in SFL.
Algorithm 2 BadSFL
@@ 4 BadSFL
Overview. To overcome the challenges, we propose BadSFL, as detailed in Algorithm 2. BadSFL mainly consists of 4 steps: Step 1: Initialization. The attacker initiates the attack by downloading the global model $w_{g}$ and controlling the variate $c$ from the server. Subsequently, the attacker updates the local model $w_{p}$ and discriminator $D$ using the downloaded global model $w_{g}$. Step 2: GAN-based Training for Data Supplementation. The attacker performs GAN training on generator $G$ and discriminator $D$. The training terminates upon the convergence of the generator, signifying its ability to generate realistic fake samples in class $C$ that do not belong to $D^{i}$ but rather originate from other clients’ datasets. Then the generator $G$ is utilized to generate a number of samples in class $c$ forming in dataset $D_{f}$. This dataset $D_{f}$ is then merged with the attacker’s original dataset $D^{i}$ to create a new dataset $D_{c}$. Step 3: Backdoor Sample Selection and Trigger Injection. With dataset $D_{c}$, the attacker selects specific samples with a characteristic feature to serve as backdoor samples. These samples are then relabeled to a target class $x$ as the backdoor target class that is different from their original labels. The attacker organizes these manipulated samples into a separate dataset $D_{b}$ and merges it with $D_{c}$ to finalize the dataset $D_{p}$ for backdoor training. Step 4: Backdoor Model Training and Optimization. The attacker proceed to trains the local model $w_{p}$ based on the dataset $D_{p}$. During the training process, the attacker follows the equation 3 to optimize the backdoor objective. Upon convergence, the backdoor model update $\Delta w_{p}$ and the corresponding control variate $\Delta c_{p}$ are obtained and can be uploaded to the server.
4.1 GAN-based Dataset Supplementation